OpenStudy (anonymous):

differentiate y=sec^4 x - 2tan^2 x. A.4sec^2 x tan ^3 x B.-4sec^2 x tan ^3 x c. 4sec^3 x tan ^2 x d.-4 sec^3 x tan^2 x

OpenStudy (agreene):

A.

OpenStudy (anonymous):

dy/dx=(sec^2)^2-2tanx^2 (1+tan^2x)^2-2tan^2x bcoz sec^2x=1+tan^2x using (a+b)^2 1+2tan^2x+tan^4x-2tan^2x 1+tan^4x finally 4tan^3xsec^2x

OpenStudy (anonymous):

sorry.........i did't get u